ten more than 6 times a number is 4 less than 4 timees the number
Answer by Maths68(1474)

Let Number = x
Six times of the x = 6x
Four times of the x=4x
6x+10=4x-4
6x-4x=-4-10
2x=-14
x=-14/2
x=-7